# Antidistillation Sampling

### v4

@@ 1 Introduction
Large language models (LLMs) trained to produce reasoning traces have achieved strong performance on math, coding, and reasoning benchmarks Jaech et al. (2024); Anthropic (2024); Guo et al. (2025). These traces, however, serve a dual purpose. They not only enhance model performance, but also enable distillation, a process by which a secondary model replicates the original model’s capabilities by training on its generations Hinton et al. (2015); Schmidhuber (1992); Gou et al. (2021). Notably, distillation can result in substantial capability gains at a fraction of the computational cost needed to train similarly performant models from scratch.
While effective and efficient, the viability of distillation poses several downsides for companies deploying frontier reasoning models. First, returning reasoning traces represents a forfeiture of intellectual property, which can allow competitors to cheaply replicate frontier capabilities. Second, the threat of distillation incentivizes limiting user access by obscuring token probabilities or truncating reasoning traces. Finally, model safety is often not preserved by distillation, which enables the generation of harmful content Sabbaghi et al. (2025); Burgess and Newman .
@@
To illustrate our empirical results, consider a reasoning model that achieves 72% accuracy on MMLU. Naively distilling this model using greedy sampling can produce a student that reaches up to 52% accuracy. If the teacher model increases its sampling temperature, its accuracy slightly decreases (e.g., by 4%), yet the distilled student’s accuracy remains largely unchanged at 52%. In contrast, using antidistillation sampling with the same 4% reduction in teacher accuracy significantly reduces the student’s capabilities, lowering its accuracy to as low as 40%. Our findings on GSM8K Cobbe et al. (2021), MATH Hendrycks et al. (2021), and MMLU Hendrycks et al. (2020) indicate that model owners can effectively limit distillation quality via antidistillation sampling.
@@ 2 Related work
Model distillation. The prominence and effectiveness of model distillation—and, more generally, model compression Buciluǎ et al. (2006); Cheng et al. (2017); Polino et al. (2018)—is rooted in a rich literature dating back to Schmidhuber (1992), and, more recently, to Hinton et al. (2015). Since these seminal works, a growing body of literature has sought to benchmark the performance of distilled models and to algorithmically maximize the effectiveness of distillation Shridhar et al. (2023); Li and Mooney (2024); Gu et al. (2023a); Goldblum et al. (2020); Romero et al. (2014). Indeed, a variety of frontier AI labs have incorporated distillation as a core technique, both to efficiently enable frontier capabilities Times (2025); DeepSeek-AI (2025) and to improve model safety via context distillation Bai et al. (2022); Snell et al. (2022); Askell et al. (2021). However, this practice constitutes a strategic vulnerability for frontier model maintainers, given the demonstrated value of these reasoning traces (see, e.g., Jaech et al. (2024); Guan et al. (2024); Lightman et al. (2023); Uesato et al. (2022)).
@@
Model security. The setting we address in this paper—where a student model is trained on data generated by a teacher model—intersects with several aspects of model security. For instance, model extraction attacks acquire weights via query-level access, whereas training data extraction attacks are designed to harvest training data Carlini et al. (2024); Nasr et al. (2023). While antidistillation sampling may offer some protection against these attacks, such analysis remains beyond our scope. More relevant is the literature on data poisoning, where maliciously crafted data is injected into a model’s training set to induce specific downstream effects (see, e.g., Goldblum et al. (2022)). In this vein, Rando and Tramèr (2024) show the effectiveness of adding backdoors to preference data, sabotaging LLMs finetuned with RLHF. Our contribution bridges data poisoning and privacy techniques to protect the valuable knowledge encoded in frontier models.
@@
Distillation prevention. A related line of work has sought to develop algorithms that prevent distillation. In the context of computer vision, Ma et al. (2021) corrupt the teacher’s logits via self-training, whereas follow-up work shows that returning only the top-$k$ logits tends to harm distillation Ma et al. (2022). Also related are watermarking algorithms, which seek to adjust model logits to detect whether a model has been distilled Zhao et al. (2023); Gu et al. (2023b); Sander et al. (2024); Kirchenbauer et al. (2023). And while this family of methods preserves teacher accuracy, their static nature—each input generally yields a deterministic logit vector—presents security vulnerabilities: a distiller can learn an inverse transformation by saving input-output pairs, and thereby fine-tune to recover uncorrupted logits. To mitigate this shortcoming, Chen et al. (2025) propose a session-dynamic defense that monitors the sensitivity of a user’s queries and perturbs the logits once a threshold is crossed. In contrast, antidistillation sampling is fully dynamic; it perturbs each token’s distribution on-the-fly using gradients from a hidden proxy model, turning generation into a moving target in a similar fashion to cryptographic stream ciphers.
@@
Language model decoding. Finally, we position antidistillation sampling within the broader framework of controlled decoding for LLMs Mudgal et al. (2023), where supplementary objectives steer the decoding process. Existing approaches in this domain include using contrastive objectives to enhance generation quality Li et al. (2022), reformulating constrained decoding as an optimization problem Ji et al. (2023), and incorporating energy-based constraints Qin et al. (2022). While related, antidistillation sampling solves a different problem: by implementing a new, distillation-aware penalization term in the decoding objective, our approach poisons generated reasoning traces to undermine the performance of models fine-tuned on these outputs.
@@ 2. For each token index $t=n,n+1,\ldots,N-1$:
$\widehat{\Delta}(\>\cdot\>|x_{1:t})\leftarrow\frac{\log p(\>\cdot\>|x_{1:t};\theta_{P}+\epsilon g)-\log p(\>\cdot\>|x_{1:t};\theta_{P})}{\epsilon}$
@@ 3.2 Preliminaries
$x_{t+1}\sim\frac{1}{Z}\exp\left(\log p(\cdot|x_{1:t};\theta)/\tau\right).$ (1)
@@ 3.3 Deriving antidistillation sampling
$x_{t+1}\sim\frac{1}{Z}\exp\left(\log p(\>\cdot\>|x_{1:t};\theta_{T})/\tau+\lambda\Delta(\>\cdot\>|x_{1:t};\theta_{P})\right),$ (5)
@@
$x_{t+1}$ $\sim\frac{1}{Z}\exp\left(\log p(\>\cdot\>|x_{1:t};\theta_{T})/\tau+\lambda\widehat{\Delta}(\>\cdot\>|x_{1:t})\right).$ (11)
@@ 4.1 Experimental setup
Architectures. To demonstrate the effectiveness of antidistillation sampling in practice, we simulate realistic distillation by instantiating distinct teacher, proxy student, and actual student models. Specifically, we use deepseek-ai/DeepSeek-R1-Distill-Qwen-7B (DeepSeek-AI, 2025) as the teacher model, Qwen/Qwen2.5-3B(Yang et al., 2024) as the proxy model, and meta-llama/Llama-3.2-3B(Grattafiori et al., 2024) as the student model (we examine other architecture configurations in §4.2).
@@
Baselines. Our primary baseline is temperature sampling, which approximates standard API endpoint behavior while providing a controlled way to degrade teacher performance. Importantly, temperature sampling ensures we fairly compare trade-offs against a straightforward baseline that—like our method—degrades teacher performance by modifying the sampling procedure. One other point of comparison to this baseline is that antidistillation sampling requires two forward passes on the proxy model for each forward pass on the teacher, independent of $\lambda$. Since we choose the proxy model to be approximately half the size of the teacher, this amounts to doubling the computation needed to sample outputs from the model compared to temperature sampling. In practice, one might choose a much smaller proxy model to reduce the overhead further.
@@
  • For $\epsilon$, we empirically verify that our finite difference approximation in eq.˜10 closely matches the JVP result in eq.˜7 (see §B.1). In practice, we find that $\epsilon=10^{-2}$ works well for BFloat16 models, which is close to the minimum in Figure 7.
@@
We use a max generation length of $1024$ for both GSM8K and MMLU and $2048$ for MATH. For antidistillation sampling, we use a temperature of $\tau=0.6$; we found that sweeping between $\tau\in[0,1]$ does not significantly impact antidistillation performance. All of our experiments are performed on nodes with 8 NVIDIA H100 GPUs and we use the transformers package (Wolf et al., 2020), the trl toolkit (von Werra et al., 2020), and the accelerate library (Gugger et al., 2022).
Distillation protocol. All distillation experiments use LoRA Hu et al. (2022) with rank $128$, $\alpha=128$, and dropout probability 0. Our optimization protocol employs a learning rate of $0.0005$, weight decay coefficient of $0.1$, and gradient clipping at norm $1.0$. Training follows a cosine learning rate schedule with warm-up over the first 10% of training, batch size $32$, for $4$ epochs. These values are the result of a systematic hyperparameter sweep using the MATH dataset to find configurations that maximize student performance gain.
@@ 4.2 Controlling the utility-distillability trade-off
Recognizing that model owners typically have limited tolerance for sacrificing utility, we explicitly focus on the high-teacher-performance regime in Figure˜9. Even in this setting, where teacher performance degradation is conservative, we observe meaningful degradation in student accuracy, underscoring the method’s practical efficacy. For example, going from 90% to 89% teacher accuracy leads to the poisoned student dropping from 65% to 56% accuracy, while temperature sampling doesn’t degrade the student’s performance at all.
@@ 4.3 Diverse configurations for antidistillation sampling
To probe the efficacy of antidistillation sampling across diverse scenarios, we conduct experiments with various teacher-student configurations and datasets. Our primary setup uses Qwen teacher and proxy student models while using a Llama model for the student. We also investigate settings where all the models (student, proxy, and teacher) belong to the same architecture family—either all from the Llama architecture family or all from the Qwen architecture family—evaluated on GSM8K. Results are provided in Figure 10.
Beyond architectural variations, we also validate our finite difference approximation by comparing it with the theoretically-motivated Jacobian-vector product (JVP) implementation (not to be confused with a vector-Jacobian product used in backpropagation). However, even in modern automatic differentiation frameworks, JVPs tend to lack support for a handful of important operations, such as SDPA. Meanwhile, JVPs are more accurately computed in Float32, but flash attention (Dao et al., 2022) only supports Float16 and BFloat16. In our implementation, we then abandon the usage of flash attention, which causes the sampling speed to decrease by around eight times (due to the memory limit, the batch size has to be decreased). Thus, for the rest of our experiments, we use finite differences for the improved convenience and efficiency (See §C for more details). As shown in Figure˜8, both approaches yield similar results in practice, confirming that our computationally efficient finite difference method provides a reliable approximation to the formal gradient-based objective. Further, with our permutation sampling baseline, we provide evidence that $\Delta$ contains useful gradient information by applying random noise with the same statistical properties as $\Delta$ (see §A).
Figure 5: Distillation loss curves show that although the student’s training loss decreases across steps, antidistillation sampling effectively poisons traces, as shown by the increase in the student’s holdout loss.
@@ Appendix A Additional Baselines
We also consider a baseline perturbation to the outputs to ensure that the computation involved in antidistillation sampling is worthwhile. This method adds random perturbations to the logits and we call this noisy sampling. While many choices of how to add noise to the output of an LLM exist, we find that matching the statistics of the perturbations computed by antidistillation sampling is the best way to find interventions that lead to the same teacher accuracy. Therefore, we randomly permute and flip the sign of the perturbations computed with antidistillation sampling to execute permutation sampling, a specific type of noisy sampling; we show the results of perturbation sampling in Figure 6.
@@
Appendix B Hyperparameters
B.1 Verifying finite difference approximation
We empirically verify that the finite difference in eq.˜10 behaves as expected by computing the relative error between the finite difference result and term produced from autograd. As shown in fig.˜7, we see it well approximates the autograd computed result for appropriately chosen step size. Here we compute $\left\langle\nabla\ell\left(\theta_{P}\right),\nabla_{\theta_{P}}\log p(x_{t+1}|x_{1:t};\theta_{P})\right\rangle$ and stack the different values of $x_{t+1}$ into a $V$ dimensional vector $\widehat{\Delta}$ and compare to the autograd vector $\Delta$. We compute relative error being sensitive only to the direction as
@@ B.1 Verifying finite difference approximation
We run this numerical experiment using Qwen/Qwen2.5-3B. Here we demonstrate that the finite difference can be used to estimate the derivatives in the low precision bfloat16 format. In particular, too small an $\epsilon$ leads to round-off error in the perturbation and too large $\epsilon$ leads to high truncation error in the Taylor expansion, with a sweet spot in the middle. The actual choice of $\epsilon$ may depends heavily on the model size (and numerical precision), so we recommend choosing this value on the exact model in question. In our actual experiment, we pick $\epsilon$ empirically to be $10^{-2}$ as suggested by fig.˜7.
@@
B.2 How We Made The Graphs
[…]
Appendix C Jacobian-Vector Product and Finite Difference
@@
There are many practical considerations that prevent us from efficiently implementing a memory-friendly JVP computations—for instance, the lack of support for popular attention kernels. At the same time, JVP is better computed in Float32 precisions.
@@
In Figure˜8, we further show that even using Float32 precision, finite difference approximation can still outperform JVP in AD sampling.

### v5

@@ 1 Introduction
Large language models (LLMs) trained to produce reasoning traces have achieved strong performance on math, coding, and reasoning benchmarks [1, 2, 3]. These traces, however, serve a dual purpose. They not only enhance model performance, but also enable distillation, a process by which a secondary model replicates the original model’s capabilities by training on its generations [4, 5, 6]. Notably, distillation can result in substantial capability gains at a fraction of the computational cost needed to train similarly performant models from scratch.
While effective and efficient, the viability of distillation poses several downsides for companies deploying frontier reasoning models. First, returning reasoning traces represents a forfeiture of intellectual property, which can allow competitors to cheaply replicate frontier capabilities. Second, the threat of distillation incentivizes limiting user access by obscuring token probabilities or truncating reasoning traces. Finally, model safety is often not preserved by distillation, which enables the generation of harmful content [7, 8].
@@
To illustrate our empirical results, consider a reasoning model that achieves 72% accuracy on MMLU. Naively distilling this model using greedy sampling can produce a student that reaches up to 52% accuracy. If the teacher model increases its sampling temperature, its accuracy slightly decreases (e.g., by 4%), yet the distilled student’s accuracy remains largely unchanged at 52%. In contrast, using antidistillation sampling with the same 4% reduction in teacher accuracy significantly reduces the student’s capabilities, lowering its accuracy to as low as 40%. Our findings on GSM8K [9], MATH [10], and MMLU [11] indicate that model owners can effectively limit distillation quality via antidistillation sampling.
@@ 2 Related work
Model distillation. The prominence and effectiveness of model distillation—and, more generally, model compression [12, 13, 14]—is rooted in a rich literature dating back to Schmidhuber [5], and, more recently, to Hinton et al. [4]. Since these seminal works, a growing body of literature has sought to benchmark the performance of distilled models and to algorithmically maximize the effectiveness of distillation [15, 16, 17, 18, 19]. Indeed, a variety of frontier AI labs have incorporated distillation as a core technique, both to efficiently enable frontier capabilities [20, 21] and to improve model safety via context distillation [22, 23, 24]. However, this practice constitutes a strategic vulnerability for frontier model maintainers, given the demonstrated value of these reasoning traces (see, e.g., [1, 25, 26, 27]).
@@
Model security. The setting we address in this paper—where a student model is trained on data generated by a teacher model—intersects with several aspects of model security. For instance, model extraction attacks acquire weights via query-level access, whereas training data extraction attacks are designed to harvest training data [28, 29]. While antidistillation sampling may offer some protection against these attacks, such analysis remains beyond our scope. More relevant is the literature on data poisoning, where maliciously crafted data is injected into a model’s training set to induce specific downstream effects (see, e.g., [30]). In this vein, Rando and Tramèr [31] show the effectiveness of adding backdoors to preference data, sabotaging LLMs finetuned with RLHF. Our contribution bridges data poisoning and privacy techniques to protect the valuable knowledge encoded in frontier models.
@@
Distillation prevention. A related line of work has sought to develop algorithms that prevent distillation. In the context of computer vision, Ma et al. [32] corrupt the teacher’s logits via self-training, whereas follow-up work shows that returning only the top-$k$ logits tends to harm distillation [33]. Also related are watermarking algorithms, which seek to adjust model logits to detect whether a model has been distilled [34, 35, 36, 37]. And while this family of methods preserves teacher accuracy, their static nature—each input generally yields a deterministic logit vector—presents security vulnerabilities: a distiller can learn an inverse transformation by saving input-output pairs, and thereby fine-tune to recover uncorrupted logits. To mitigate this shortcoming, Chen et al. [38] propose a session-dynamic defense that monitors the sensitivity of a user’s queries and perturbs the logits once a threshold is crossed. In contrast, antidistillation sampling is fully dynamic; it perturbs each token’s distribution on-the-fly using gradients from a hidden proxy model, turning generation into a moving target in a similar fashion to cryptographic stream ciphers.
@@
Language model decoding. Finally, we position antidistillation sampling within the broader framework of controlled decoding for LLMs [39], where supplementary objectives steer the decoding process. Existing approaches in this domain include using contrastive objectives to enhance generation quality [40], reformulating constrained decoding as an optimization problem [41], and incorporating energy-based constraints [42]. While related, antidistillation sampling solves a different problem: by implementing a new, distillation-aware penalization term in the decoding objective, our approach poisons generated reasoning traces to undermine the performance of models fine-tuned on these outputs.
@@ 2. For each token index $t=n,n+1,\ldots,N-1$:
$\widehat{\Delta}(\>\cdot\>|x_{1:t})\leftarrow\frac{\log p(\>\cdot\>|x_{1:t};\theta_{P}+\epsilon g)-\log p(\>\cdot\>|x_{1:t};\theta_{P}-\epsilon g)}{2\epsilon}$
@@ 3.2 Preliminaries
$x_{t+1}\sim\frac{1}{Z}\exp\left(\frac{1}{\tau}\log p(\cdot|x_{1:t};\theta)\right).$ (1)
@@ 3.3 Deriving antidistillation sampling
$x_{t+1}\sim\frac{1}{Z}\exp\left(\frac{1}{\tau}\log p(\>\cdot\>|x_{1:t};\theta_{T})+\lambda\Delta(\>\cdot\>|x_{1:t};\theta_{P})\right),$ (5)
@@
$x_{t+1}$ $\sim\frac{1}{Z}\exp\left(\frac{1}{\tau}\log p(\>\cdot\>|x_{1:t};\theta_{T})+\lambda\widehat{\Delta}(\>\cdot\>|x_{1:t})\right).$ (11)
@@ 4.1 Experimental setup
Architectures. To demonstrate the effectiveness of antidistillation sampling in practice, we simulate realistic distillation by instantiating distinct teacher, proxy student, and actual student models. Specifically, we use deepseek-ai/DeepSeek-R1-Distill-Qwen-7B [21] as the teacher model, Qwen/Qwen2.5-3B [43] as the proxy model, and meta-llama/Llama-3.2-3B [44] as the student model (we examine other architecture configurations in §4.3).
@@
Baselines. Our primary baseline is temperature sampling, which approximates standard API endpoint behavior while providing a controlled way to degrade teacher performance. Importantly, temperature sampling ensures we fairly compare trade-offs against a straightforward baseline that—like our method—degrades teacher performance by modifying the sampling procedure. One other point of comparison to this baseline is that antidistillation sampling requires two forward passes on the proxy model for each forward pass on the teacher, independent of $\lambda$. Since we choose the proxy model to be approximately half the size of the teacher, this amounts to doubling the computation needed to sample outputs from the model compared to temperature sampling. In practice, one might choose a much smaller proxy model to reduce the overhead further (see Figure 11). We also introduce a permutation sampling baseline in §A that preserves the statistical properties of $\widehat{\Delta}$ while scrambling the gradient information, providing definitive evidence that the computational overhead of producing $\widehat{\Delta}$ is necessary to achieve our non-distillability objective.
@@
  • For $\epsilon$, we empirically verify that our finite difference approximation in eq.˜10 closely matches the JVP result in eq.˜7 (see §B). In practice, we find that $\epsilon=10^{-2}$ works well for BFloat16 models, which is close to the minimum in Figure˜7.
@@
We use a max generation length of $1024$ for both GSM8K and MMLU and $2048$ for MATH. For antidistillation sampling, we use a temperature of $\tau=0.6$; we found that sweeping between $\tau\in[0,1]$ does not significantly impact antidistillation performance. All of our experiments are performed on nodes with 8 NVIDIA H100 GPUs and we use the transformers package [46], the trl toolkit [47], and the accelerate library [48].
Distillation protocol. All distillation experiments use LoRA [49] with rank $128$, $\alpha=128$, and dropout probability 0 Our optimization protocol employs a learning rate of $0.0005$, weight decay coefficient of $0.1$, and gradient clipping at norm $1.0$. Training follows a cosine learning rate schedule with warm-up over the first 10% of training, batch size $32$, for $4$ epochs. These values are the result of a systematic hyperparameter sweep using the MATH dataset to find configurations that maximize student performance gain.
@@ 4.2 Controlling the utility-distillability trade-off
Recognizing that model owners typically have limited tolerance for sacrificing utility, we explicitly focus on the high-teacher-performance regime in Figure˜9. Even in this setting, where teacher performance degradation is conservative, we observe meaningful degradation in student accuracy, underscoring the method’s practical efficacy. For example, going from 90% to 89% teacher accuracy leads to the poisoned student dropping from 65% to 56% accuracy, while temperature sampling doesn’t degrade the student’s performance at all. For illustrative examples comparing traces generated across comparable $\lambda$ and $\tau$ settings, see §E.
@@ 4.3 Diverse configurations for antidistillation sampling
To probe the efficacy of antidistillation sampling across diverse scenarios, we conduct experiments with various teacher-student configurations and datasets. Our primary setup uses Qwen teacher and proxy student models while using a Llama model for the student. We also investigate settings where all the models (student, proxy, and teacher) belong to the same architecture family—either all from the Llama architecture family or all from the Qwen architecture family—evaluated on GSM8K. Results are provided in Figure˜10.
Beyond architectural variations, we also validate our finite difference approximation by comparing it with the theoretically-motivated Jacobian-vector product (JVP) implementation (not to be confused with a vector-Jacobian product used in backpropagation) (see §C). Both approaches yield similar results in practice, confirming that our computationally efficient finite difference method provides a reliable approximation to the formal gradient-based objective.
Figure 5: Distillation loss curves show that although the student’s training loss decreases across steps, antidistillation sampling effectively poisons traces, as shown by the increasing student’s holdout loss.
@@ 4.4 Generalizing from the proxy model
In all our experiments, we keep the size discrepancy between proxy and student model relatively small (both are 3B models). However, Figure˜11 demonstrates that our method remains effective even when proxy and student models differ in size, suggesting robustness to this architectural mismatch.
@@ Appendix A Additional Baselines
We also consider a baseline perturbation to the outputs to ensure that the computation involved in antidistillation sampling is worthwhile. This method adds random perturbations to the logits and we call this noisy sampling. While many choices of how to add noise to the output of an LLM exist, we find that matching the statistics of the perturbations computed by antidistillation sampling is the best way to find interventions that lead to the same teacher accuracy. Therefore, we randomly permute and flip the sign of the perturbations computed with antidistillation sampling to execute permutation sampling, a specific type of noisy sampling; we show the results of perturbation sampling in Figure˜6.
@@
Appendix B Verifying hyperparameter choice $\epsilon$
We empirically verify that the finite difference in eq.˜10 behaves as expected by computing the relative error between the finite difference result and term produced from autograd. As shown in Figure˜7, we see it well approximates the autograd computed result for appropriately chosen step size. Here we compute $\left\langle\nabla\ell\left(\theta_{P}\right),\nabla_{\theta_{P}}\log p(x_{t+1}|x_{1:t};\theta_{P})\right\rangle$ and stack the different values of $x_{t+1}$ into a $V$ dimensional vector $\widehat{\Delta}$ and compare to the autograd vector $\Delta$. We compute relative error being sensitive only to the direction as
@@ Appendix B Verifying hyperparameter choice $\epsilon$
We run this numerical experiment using Qwen/Qwen2.5-3B. Here we demonstrate that the finite difference can be used to estimate the derivatives in the low precision bfloat16 format. In particular, too small an $\epsilon$ leads to round-off error in the perturbation and too large $\epsilon$ leads to high truncation error in the Taylor expansion, with a sweet spot in the middle. The actual choice of $\epsilon$ may depends heavily on the model size (and numerical precision), so we recommend choosing this value on the exact model in question. In our actual experiment, we pick $\epsilon$ empirically to be $10^{-2}$ as suggested by Figure˜7.
@@
Appendix C Verifying finite difference approximation
@@
Even in modern automatic differentiation frameworks, There are many practical considerations that prevent us from efficiently implementing memory-friendly JVP computations. JVPs tend to lack support for a handful of important operations, such as SDPA. While JVPs are more accurately computed in Float32 flash attention [50] only supports Float16 and BFloat16. In our implementation of JVPs, we then abandon the usage of flash attention, which causes the sampling speed to decrease by around eight times (due to the memory limit, the batch size has to be decreased).
@@
In Figure˜8, we further show that even using Float32 precision, finite difference approximation can still outperform JVP in AD sampling. Thus, for the rest of our experiments, we use finite differences for the improved convenience and efficiency.
Appendix D How We Made The Graphs
[…]